OpenStudy (anonymous):

Consider curve -x^2 + 2xy + 2y^2 = 3 Show that (x +2y)y' - x + y =0 ?

OpenStudy (raden):

derive implcitly from both sides : -x^2 + 2xy + 2y^2 = 3 -2x + 2y + 2x y' + 4y y' = 0 divide by 2 from both sides, will giving us -x + y + x y' + 2y y' = 0 y' (x+2y) - x + y = 0 QED